OpenStudy (hannahc234):

Solve the equation by completing the square. Round to the nearest tenth. x^2+8x=10 A. 1.1, 9.1 B. 1.1, -9.1 C. -1.1, 9.1 D. -1.1, -9.1

OpenStudy (anonymous):

what is half of 8?

OpenStudy (hannahc234):

4

OpenStudy (anonymous):

ok good, and what is \(4^2\)?

OpenStudy (hannahc234):

16

OpenStudy (anonymous):

so go right from \[x^2+8x=10\] so \[(x+4)^2=10+16\\ (x+4)^2=26\]

OpenStudy (anonymous):

You're trying to find (x+?)^2 that fulfills the requirement of having 8x this means that the ? is 8/2=4 So now because the ? is 4, and 4^2=16, you add 16 to both sides giving you: (x+4)^2=10+16 Square root both sides: \[x+4=\sqrt26\] subtract 4 from both sides and you have: \[x=\sqrt26 - 4\] x=5.10-4 x=-5.10-4 x=1.1 and x=-9.1

OpenStudy (anonymous):

take the square root of both sides, don't forget the \(\pm\) get \[x+4=\pm\sqrt{26}\] then a calculator

OpenStudy (hannahc234):

Thank you guys so much!